Angle ABD=(12x-12), Angle BDE=(3x+48). Both are supplementary. Find the measures of both angles.
Answer by cleomenius(959)

12x - 12 + 3x + 48 = 180
15x + 36 = 180
15x = 144
x = 9.6
12(9.6) - 12 = 103.2 degrees.
3(9.6) 28.8 + 48 = 76.8 degrees.
103.2 + 76.8 = 180 degrees.
This does check.
Cleomenius.